# How Numerical Precision Affects Mathematical Reasoning Capabilities of LLMs

### v1

@@ 5 Standard-Precision Transformers Are Sufficient for Arithmetic Tasks
We first establish that, under logarithmic precision, a Transformer with constant depth and dimension can solve both the integer addition and iterated addition tasks for arbitrarily large input lengths, as shown in Theorems 5.1 and 5.2.
@@
For any integers $n$ and $p$, there exists a logarithmic-precision Transformer with constant depth $L$ and constant hidden dimension $d$ (independent of $n$) that can generate the correct output for any input on the $\operatorname{ADD}(n,p)$ task.
@@
For any integers $n$, $k$, and $p$, there exists a logarithmic-precision Transformer with constant depth $L$ and constant hidden dimension $d$ (independent of $n$ and $k$) that can generate the correct output for any input on the $\operatorname{IterADD}(n,k,p)$ task.
We now turn to integer multiplication. As established in Theorem 5.3, a logarithmic-precision Transformer with constant depth and polynomial hidden dimensions is capable of solving the integer multiplication task.
@@
For any integers $n$, $l$, and $p$, there exists a logarithmic-precision Transformer with constant depth and hidden dimensions $O(n^{2})$ that can generate the correct output for any input on the $\operatorname{MUL}(n,l,p)$ task.
@@
The Reason for Increased Expressiveness. In Section 4, we highlighted low-precision Transformers’ difficulties in performing elementary arithmetic operations. A critical insight is that low-precision Transformers struggle to store intermediate results, even for simple operations like adding two digits. This issue is mitigated by using standard precision. For instance, float32 can accurately represent integers within a sufficiently large range to store intermediate results during computation. Thus, standard precision allows Transformers to perform arithmetic tasks more effectively.
@@
Figure 2: Model performance on different tasks in base-2. Within each sub-figure, the x-axis represents the maximum digits length and the y-axis represents the accuracy gained by each model. The figure indicates that, for all tasks, Transformers utilizing float32 (32-bit precision) with 3 layers and 5 layers outperform their bfloat16 (16-bit precision) counterparts.
@@ 6 Experiments
In the preceding sections, we employed complexity theory to demonstrate that low-precision Transformers face significant challenges in performing elementary arithmetic tasks. To validate these theoretical insights, we conduct a series of experiments to compare the performance of Transformer models under different precisions. The results provide empirical evidence that the model’s ability to execute arithmetic operations drops as precision decreases, reinforcing our theoretical results.
@@ 6.1 Experimental Setup
Tasks and datasets.
We evaluate three elementary arithmetic tasks: integer addition, iterated addition, and integer multiplication, as presented in Figure 1. Each task involves a series of experiments with base $p=2,10$ and varying choices of digit length $n$. For integer addition, we examine the addition of integers in both base-2 and base-10, with digit lengths $n\in\{4,8,16,32,64\}$. For iterated addition, we examine the addition of three numbers in base-2, with digit lengths $n\in[2,11]$, as well as in base-10, with digit lengths $n\in[1,4]$. Similarly, for integer multiplication, we run experiments in base-2 with digit lengths $n\in[2,14]$, and in base-10 with digit length $n\in[2,5]$. We dynamically generated datasets for both training and testing, with further details on dataset construction available in Appendix F.
Model configurations.
For all experiments, we use Transformer models with hidden dimension $d=256$, heads $H=4$, and model depth $L\in\{3,5\}$. The causal self-attention layer employs rotary positional embeddings Su et al. (2024), which replace the traditional sinusoidal embeddings. For activation, we choose NewGeLU, the variant of GeLU Hendrycks and Gimpel (2016), and apply Xiaver initialization across the parameters Glorot and Bengio (2010).
Figure 3: We evaluate model performance on iterated addition tasks involving three numbers and integer multiplication tasks. Each sub-figure presents a comparison of the performance between float32 (32-bit precision) and bfloat16 (16-bit precision).
Model training and inference.
[…]
@@ 6.2 Experimental Results
Integer addition proves to be too easy, maintaining more than 94% accuracy even as digit length increases to 32 in both base-2 and base-10 with both float32 and bfloat16 (detailed results in Appendix F). The results of iterated addition and multiplication in base-2 are shown in Figure 2, while the corresponding results in base-10 are shown in Figure 3. Each sub-figure corresponds to a task with the x-axis representing the maximum length of the addends and the multiplicands respectively, and the y-axis representing the test accuracy. In the context of the iterated addition task, the accuracy shows a significantly more pronounced decline in 16-bit precision compared to 32-bit precision as the maximum digits length increases, across all model depths. Specifically, the 16-bit precision exhibits a marked decline in performance for lengths ranging from 7 to 10 in base-2, whereas the 32-bit precision maintains near-complete accuracy across these ranges. In base-10, 32-bit can achieve $90\%$ correctness in contrast to the 16-bit representation, which struggles to yield accurate results. In the context of multiplication tasks, as the maximum digit length increases, 32-bit precision achieves significantly higher accuracy compared to 16-bit precision. Notably, when the length reaches 13 in base-2, the accuracy of 16-bit drops to a very low level, indicating their inability to generate correct results. Additionally, in base-10, we observe a noticeable reduction in accuracy when transitioning from 32-bit to 16-bit precision, particularly for inputs with a maximum length of 3 in 3-layer models and a maximum length of 4 in 5-layer models. This suggests that the differences of precision for performing elementary arithmetic tasks are consistent with our theoretical results.
@@ 6.3 Further Experiments on LLMs
Figure 4: The performance of LLAMA 3.1 8B Instruct on various tasks in base-10. Within each sub-figure, we compare the performance of bfloat16 (16-bit precision) and int4 (4-bit precision).
We further conducted an extensive set of experiments on LLMs to empirically support our theoretical findings. Specifically, we evaluated the models’ performance across three elementary arithmetic tasks, as illustrated in Figure 1. In the integer addition task, we tested the addition of two base-10 integers, varying their digit lengths from $1$ to $13$. For the iterated addition task, we expanded the setup to include the addition of three and five base-10 numbers, also with digit lengths ranging from $1$ to $9$ and $1$ to $5$. Similarly, in the integer multiplication task, we examined the multiplication of two base-10 numbers, with digit lengths spanning from $1$ to $5$. The process of generating data is the same as in previous experiments.
@@
All experiments utilized the LLaMA 3.1 8B Instruct model (Dubey et al., 2024) as the base model, employing a few-shot learning approach for inference. Detailed specifications of the prompt construction and generation parameters can be found in the Appendix F. During inference, the LLMs were tasked with producing exact solutions to the given arithmetic problems. For each task, we evaluate the model on 1k samples to compute the accuracy serving as the evaluation metric. Since the LLaMA 3.1 model is trained with bfloat16 precision, we adopted its performance under bfloat16 as our baseline. To further investigate the impact of reduced numerical precision, we additionally evaluated the model’s performance when quantized to 4-bit precision via the AWQ algorithm (Lin et al., 2024).
@@
The results of the experiments are shown in Figure 4. Each sub-figure presents the results of a task, where the x-axis denotes the maximum length of the addends or multiplicands, and the y-axis represents the test accuracy. In the integer addition task, reducing the numerical precision from 16-bit to 4-bit has no observable effect on accuracy when the number length is less than or equal to 7, and only a slight reduction in accuracy is observed when the number length exceeds 7. This comparable performance between 16-bit and 4-bit precision supports our theorem, which asserts that both low-precision and standard-precision arithmetic can successfully handle integer addition tasks. In contrast, for iterated addition and integer multiplication tasks, the reduction in numerical precision results in a significant decrease in accuracy. Specifically, in the iterated addition task, accuracy drops by nearly 50% as the length of the addends increases. The impact is even more pronounced in the integer multiplication task, where performance deteriorates further. These experimental findings support our theoretical results that numerical precision is a critical factor in the success of iterated addition and integer multiplication tasks. Overall, the results underscore the consistency between the precision requirements for these elementary arithmetic tasks and our theoretical predictions.
@@ 7 Conclusion
In this work, we have theoretically analyzed the impact of numerical precision on LLMs for mathematical reasoning. By focusing on three elementary arithmetic tasks, integer addition, iterated addition, and integer multiplication, we constructively demonstrate that the Transformers operating under standard precision can handle these tasks effectively. In contrast, Transformers with low precision struggle with complex arithmetic tasks, excelling only at integer addition. Extensive experimental results corroborate our theoretical findings, showing that standard precision models outperform low precision ones. We believe this study offers valuable insights for the development of more powerful LLMs in mathematical reasoning.
@@ A.2 Computational Powers of Transformers
Chain-of-Thought (CoT) prompting (Wei et al., 2022b) plays a crucial role in tasks requiring complex reasoning structures, and several studies aim to understand its underlying mechanisms. For instance, Feng et al. (2023); Li et al. (2024) analyzed CoT from an expressiveness perspective, and Yang et al. (2024); Wen et al. (2024) examined CoT across more different model variants. In-context learning Brown et al. (2020); Garg et al. (2022) is another powerful aspect of LLMs. Some theoretical work has shown that in-context learning can be explained through gradient descent Akyürek et al. (2022); Dai et al. (2023); Von Oswald et al. (2023), while others attribute it to the induction heads mechanism Elhage et al. (2021); Olsson et al. (2022).
@@ B.2 Constant-precision Transformer
For any $x\in{\mathbb{R}}$, its representation under this floating-point format is determined by rounding to the nearest value in ${\mathbb{F}}$. In the event of a tie, we select the number with the smaller absolute value. In this paper, we focus on the case where $e=0$, which means all representable numbers take the form $S\cdot 2^{-s}$, with $S\in{\mathbb{Z}}$ such that $-2^{-2s}+1\leq S\leq 2^{2s}-1$. However, this is necessary only for Theorem 4.1, while Theorems 4.2 and 4.3 do not depend on specific numerical representations.
Li et al. (2024) demonstrated that constant-depth Transformers with constant precision belong to the complexity class $\mathsf{AC}^{0}$.
@@ B.3 Logarithmic-precision Transformer
A key limitation of constant-precision representation is that it fails to capture the input size $n$ within a single neuron. To address this, we consider logarithmic precision, allowing for $O(\log n)$ bits for numerical representations. Logarithmic-precision Transformers possess several advantageous properties Feng et al. (2023); Feng and Zhong (2023):

### v2

@@ 5 Standard-Precision Transformers Are Sufficient for Arithmetic Tasks
We first establish that, under logarithmic precision, a Transformer with constant depth and dimension can solve both the integer addition and iterated addition tasks for arbitrarily large input lengths, as shown in Theorems 5.1 and 5.2. The detailed proof of Theorems 5.1 and 5.2 is presented in Sections E.1 and E.2.
@@
Fix integers $p\geq 2$ and $c\in\mathbb{N}^{*}$. Consider the tokenizer ${\bm{T}}_{c}$ defined in Eq. 1 for processing the input and output sequences. There exists a logarithmic-precision Transformer with constant depth and hidden dimension (independent of $n$) that can generate the correct output for any input on the $\operatorname{ADD}_{p}(n)$ task.
@@
Fix integers $p\geq 2$ and $c\in\mathbb{N}^{*}$. Consider the tokenizer ${\bm{T}}_{c}$ defined in Eq. 1 for processing the input and output sequences. For any integers $n$ and $k$, there exists a logarithmic-precision Transformer with constant depth and hidden dimension $d$ (independent of $n$ and $k$) that can generate the correct output for any input on the $\operatorname{IterADD}_{p}(n,k)$ task.
We now turn to integer multiplication. As established in Theorem 5.3, a logarithmic-precision Transformer with constant depth and polynomial hidden dimensions is capable of solving the integer multiplication task. The detailed proof of Theorem 5.3 is presented in Section E.3.
@@
Fix integers $p\geq 2$ and $c\in\mathbb{N}^{*}$. Consider the tokenizer ${\bm{T}}_{c}$ defined in Eq. 1 for processing the input and output sequences. For any integers $n$ and $l\leq 2n$, there exists a logarithmic-precision Transformer with constant depth (independent of $n$ and $k$) and hidden dimensions $O(n^{2})$ that can generate the correct output for any input on the $\operatorname{MUL}_{p}(n,l)$ task.
@@
The Reason for Increased Expressiveness. The transition from constant precision to logarithmic precision enables Transformers to process and represent large numbers effectively, thereby expanding their expressiveness beyond the capabilities of low-precision models. In particular, the expressiveness of a logarithmic-precision Transformer with polynomial size and bounded depth is upper-bounded by the computational complexity class $\mathsf{TC}^{0}$ (Merrill and Sabharwal, 2023). Leveraging this increased precision, we constructively prove that logarithmic-precision Transformers are sufficient for solving these arithmetic tasks. These results underscore the critical role of numerical precision in enhancing the expressiveness of Transformer architectures.
@@
Figure 2: Model performance on different tasks in base-2. Within each sub-figure, the x-axis represents the maximum digits length and the y-axis represents the accuracy gained by each model. The figure indicates that, for all tasks, Transformers utilizing float32 with 3 layers and 5 layers outperform their bfloat16 counterparts.
@@ 6 Experiments
In the preceding sections, we employ complexity theory to demonstrate that low-precision Transformers face significant challenges in performing elementary arithmetic tasks. To validate these theoretical insights, we conduct a series of experiments to compare the performance of Transformers under different precisions. The results provide empirical evidence that the model’s ability to execute arithmetic operations drops as precision decreases, reinforcing our theoretical results.
@@ 6.1 Experimental Setup
Tasks and datasets. We evaluate three elementary arithmetic tasks: integer addition, iterated addition, and integer multiplication, as presented in Figure 1. Each task involves a series of experiments with base $p=2,10$ and varying choices of digit length $n$. For integer addition, we examine the addition of integers in both base-2 and base-10, with digit lengths $n\in\{4,8,16,32,64\}$. For iterated addition, we examine the addition of three numbers in base-2, with digit lengths $n\in[2,11]$, as well as in base-10, with digit lengths $n\in[1,4]$. Similarly, for integer multiplication, we run experiments in base-2 with digit lengths $n\in[2,14]$, and in base-10 with digit length $n\in[2,5]$. Both training data and test data are dynamically generated. We use a batch size of 512 with 100k steps, resulting in a total training dataset size of 51.2M. Further details regarding the data generation function and the construction of datasets are provided in Algorithms 4 and 5.
Training and Evaluation. All experiments use Transformers as the backbone. We trained models with 3 and 5 layers and evaluated their performance on each task. Detailed model and training configurations are listed in Tables 2 and 3. No prompts or chat templates were added to the dataset. The models were trained with cross-entropy loss over the answer tokens. During evaluation, the models were required to produce exact answers, with accuracy reported as the evaluation metric. For each task, accuracy was computed over 50k test samples. To assess the impact of numerical precision, experiments were conducted with float32 and bfloat16.
Figure 3: Model performance on iterated addition tasks involving three numbers and integer multiplication tasks. Each sub-figure presents a comparison of the performance between float32 and bfloat16.
@@ 6.2 Experimental Results
Integer addition proved relatively simple, maintaining over 94% accuracy even as digit lengths increased to 32 across both base-2 and base-10 for both float32 and bfloat16 (see Section F.3).
The results for iterated addition and multiplication in base-2 are shown in Figure 2, while the corresponding base-10 results are presented in Figure 3. In each sub-figure, the x-axis represents the maximum digit length for addends or multiplicands, while the y-axis indicates test accuracy.
For iterated addition, accuracy under bfloat16 declined significantly as the digit length increased, while float32 consistently achieved near-perfect accuracy across all model depths. Specifically, in base-2, 16-bit precision exhibited a pronounced decline for digit lengths between 7 and 10, whereas 32-bit precision maintained high accuracy. In base-10, at digit lengths up to 10, float32 achieved over 90% accuracy, whereas bfloat16 struggled to produce correct results.
In the multiplication task, the gap between the two precisions became even more apparent as digit lengths increased. For example, at a digit length of 13 in base-2, 16-bit precision accuracy dropped sharply, signifying its inability to handle such inputs. Similarly, in base-10, 16-bit precision showed a marked reduction in accuracy, particularly for inputs with lengths of 3 in 3-layer models and lengths of 4 in 5-layer models. These results underscore the critical role of precision in achieving reliable performance for elementary arithmetic tasks, consistent with our theoretical findings.
@@ 6.3 Further Experiments on LLMs
Figure 4: The performance of LLAMA-3.1-8B Instruct model on arithmetic tasks in base-10. In each sub-figure, we compare the original model in bfloat16 and the quantized model in int4, alongside fine-tuned models, with LoRA using bfloat16 and QLoRA using int4.
To further substantiate our theoretical results, we conducted additional experiments on LLMs, specifically evaluating the LLAMA-3.1-8B Instruct model on elementary arithmetic tasks.
Task Description. For integer addition, we tested the addition of two base-10 integers with digit lengths ranging from $1$ to $13$. For iterated addition, we extended the task to include three and five base-10 numbers, with digit lengths spanning $1$ to $9$ and $1$ to $5$, respectively. For integer multiplication, we evaluated the multiplication of two base-10 numbers, with digit lengths varying from $1$ to $5$. Data generation followed the same procedure as earlier experiments, with details provided in Algorithms 4 and 5.
Model Configuration. All experiments used the LLAMA-3.1-8B Instruct model (Dubey et al., 2024). To study the effects of reduced precision, we evaluated the model under four settings:
  • Original model operating under bfloat16
  • Quantized model operating under int4
  • Fine-tuned model using LoRA (bfloat16)
  • Fine-tuned model using QLoRA (int4)
The baseline configuration employs the original LLaMA-3.1-8B Instruct model, which operates under bfloat16 precision. To assess the effects of reduced precision, we applied 4-bit quantization using the AWQ algorithm (Lin et al., 2024). Further, we fine-tuned the model using LoRA and QLoRA (Hu et al., 2021; Dettmers et al., 2024). The fine-tuning configurations for LoRA and QLoRA are listed in Table 6 in Section F.4. For the LoRA fine-tuning experiments, model weights were maintained in bfloat16. In contrast, the QLoRA experiments extended this setup by enabling 4-bit quantization, represented by the int4. Fine-tuning was performed individually for each task. Furthermore, we add a baseline of GPT-4o (OpenAI, 2023) as a reference, whose results are listed in Table 9.
Dataset for Fine-tuning. We generated the fine-tuning data for both multiplication and addition tasks, with both multiplicands and addends varying in length from 1 to 9. The dataset comprised a total of 60k samples, including 5k samples with lengths between 1 and 6, and 10k samples with lengths between 7 and 9. The generation process of the dataset is the same as in previous experiments. Furthermore, we add the few-shot learning prompt to the raw dataset and apply the LLaMA chat template for data preprocessing. The prompt for few-shot learning can be found in Tables 7 and 8.
@@
Evaluation. For the evaluation, we employ a few-shot learning approach for inference. The prompts are the same as the prompts of fine-tuning dataset and can be found in Section F.1. The generation configurations for LLMs can be also found in Table 5 in Section F.4. During inference, the LLMs were tasked with producing exact solutions to the given arithmetic problems. Both the original model and the model fine-tuned with LoRA are evaluated using bfloat16, whereas the quantized model and the model fine-tuned with QLoRA are evaluated using int4. For each task, we evaluate the model on 1k samples to compute the accuracy serving as the evaluation metric.
@@
The results of the experiments are shown in Figure 4. Each sub-figure presents the results of a task, where the x-axis denotes the maximum length of the addends or multiplicands, and the y-axis represents the test accuracy. For each task, reducing numerical precision in both the original and fine-tuned models leads to a significant decrease in accuracy. Specifically, in the iterated addition task for 3 numbers, accuracy drops by nearly $20\%$ as the length of the addends increases. Similarly, for models fine-tuned with QLoRA and LoRA, lowering precision also results in a decline in accuracy. Furthermore, in some cases, even after fine-tuning a low-precision model with QLoRA, the performance does not surpass that of the original model with standard precision. These experimental findings support our theoretical results that numerical precision is a critical factor in the success of iterated addition and integer multiplication tasks. Overall, the results underscore the consistency between the precision requirements for these elementary arithmetic tasks and our theoretical predictions.
@@ 7 Conclusion
In this work, we have theoretically analyzed the impact of numerical precision on LLMs for arithmetical reasoning. By focusing on three elementary arithmetic tasks, integer addition, iterated addition, and integer multiplication, we demonstrate that the Transformers operating under standard precision can handle these tasks effectively. In contrast, Transformers with low precision struggle with complex arithmetic tasks, excelling only at integer addition. Extensive experimental results corroborate our theoretical findings, showing that standard precision models outperform low precision ones. We believe this study offers valuable insights for developing more powerful LLMs in mathematics.
@@ 8 Limitations
Acknowledgments
Liwei Wang is supported by National Science Foundation of China (NSFC62276005). Di He is supported by National Science Foundation of China (NSFC62376007). This work is partially supported by the Shanghai Committee of Science and Technology (Grant No. 21DZ1100100).
@@ A.2 Computational Powers of Transformers
Chain-of-Thought prompting (Wei et al., 2022b) plays a crucial role in tasks requiring complex reasoning structures, and several studies aim to understand its underlying mechanisms. For instance, Feng et al. (2023); Li et al. (2024) analyzed CoT from an expressiveness perspective, and Yang et al. (2024); Wen et al. (2024) examined CoT across more different model variants. In-context learning Brown et al. (2020); Garg et al. (2022) is another powerful aspect of LLMs. Some theoretical work has shown that in-context learning can be explained through gradient descent Akyürek et al. (2022); Dai et al. (2023); Von Oswald et al. (2023), while others attribute it to the induction heads mechanism Elhage et al. (2021); Olsson et al. (2022).
A.3 Scaling Laws of Precision
Concurrent works Kumar et al. (2024); Ouyang et al. (2024) explore the impact of numerical precision on scaling laws, particularly in the contexts of training and quantization. Kumar et al. (2024) introduced “precision-aware” scaling laws, demonstrating that low-precision training effectively reduces a model’s “effective parameter count” but may still be compute-optimal for larger models. Their framework unifies the effects of both training and post-training quantization. Ouyang et al. (2024) examined quantization-induced degradation (QiD), showing that larger or undertrained models exhibit greater robustness to low-bit quantization, whereas fully trained models experience significant performance degradation. While both studies underscore precision as a critical dimension in scaling laws, they leave theoretical gaps in understanding the role of precision for LLMs. Our work focuses on addressing these gaps by analyzing the impact of numerical precision on elementary arithmetic reasoning tasks.
@@ B.2 Constant-precision Transformer
For any $x\in{\mathbb{R}}$, its representation under this floating-point format is determined by rounding to the nearest value in ${\mathbb{F}}$. In the event of a tie, we select the number with the smaller absolute value.
In this paper, we focus on the case where $e=0$, which means all representable numbers take the form $S\cdot 2^{-s}$, with $S\in{\mathbb{Z}}$ such that $-2^{-2s}+1\leq S\leq 2^{2s}-1$. However, this is necessary only for Theorem 4.1, while Theorems 4.2 and 4.3 do not depend on specific numerical representations.
Definition B.2 (Constant-Precision Transformer).
A constant-precision Transformer is a Transformer in which each neuron and activation are restricted to using a constant number of bits for computation.
Li et al. (2024) demonstrated that constant-precision Transformers with constant depth belong to the complexity class $\mathsf{AC}^{0}$.
@@ B.3 Logarithmic-precision Transformer
A key limitation of constant-precision representation is that it fails to capture the input size $n$ within a single neuron. To address this, we consider logarithmic precision, allowing for $O(\log n)$ bits for numerical representations.
Definition B.3 (Logarithmic-Precision Transformer).
A logarithmic-precision Transformer is a Transformer in which each neuron and activation are allowed to use $O(\log n)$ bits for computation, where $n$ denotes the size of the input.
Logarithmic-precision Transformers possess several advantageous properties Feng et al. (2023); Feng and Zhong (2023):
@@
B.4 Tokenization Scheme
In this section, we formalize the tokenization scheme adopted in this paper and provide the necessary definitions and examples to establish a foundation for the subsequent analysis.
Definition B.4 (Tokenizer ${\bm{T}}_{c}$).
Let $\mathbf{x}=(x_{n-1}\cdots x_{0})_{p}$ denote an $n$-digit integer in base $p$. The tokenizer ${\bm{T}}_{c}$ maps $\mathbf{x}$ into $k=\lceil\frac{n}{c}\rceil$ tokens, represented as $\bm{t}=[t_{k-1},\cdots,t_{0}]$, where
$t_{i}=\begin{cases}[x_{ic},x_{ic+1},\cdots,x_{ic+c-1}],&i<k;\\ [x_{ic},x_{ic+1},\cdots,x_{n-1}],&i=k.\end{cases}$
Furthermore, for any operator (e.g., “$+$”, “$\times$”, “$=$”), the tokenizer ${\bm{T}}_{c}$ assigns each operator a single token.
Example B.5.
Consider the $5$-digit integer $13215$. Under the tokenizer ${\bm{T}}_{3}$, it is tokenized into $[13,215]$.
A key property of this tokenization scheme is that, for any fixed base $p$ and tokenizer ${\bm{T}}_{c}$, the resulting tokenized sequence can be reinterpreted as an arithmetic expression in base $p^{c}$. Specifically, there exists a one-to-one mapping $\tau$ between the vocabulary of the base-$p$ tokenizer ${\bm{T}}_{c}$ and the vocabulary of the base-$p^{c}$ tokenizer ${\bm{T}}_{1}$, such that
$\tau({\bm{T}}_{c}([a_{c-1},\cdots,a_{0}]_{p}))={\bm{T}}_{1}\left(\left[\sum_{i\in[c]}a_{i}p^{i}\right]\right).$
Proposition B.6.
Let ${\bm{a}}$ be an integer. If ${\bm{t}}={\bm{T}}_{c}({\bm{a}})=[t_{k-1},\cdots,t_{0}]$ and ${\bm{t}}^{\prime}={\bm{T}}_{1}({\bm{a}})=[t_{k-1}^{\prime},\cdots,t_{0}^{\prime}]$, then for all $i$, we have $\tau(t_{i})=t_{i}^{\prime}$.
This property is particularly significant as it allows us to abstract away the specific effects of the tokenizer ${\bm{T}}_{c}$ and focus exclusively on the case where the tokenizer is ${\bm{T}}_{1}$. This simplification is leveraged in proving the main theorems presented in this paper.
Example B.7 illustrates how a tokenized sequence in base-10, generated using the tokenizer ${\bm{T}}_{3}$, can be equivalently interpreted as a sequence in base-1000.
Example B.7.
Consider the base-10 arithmetic expression $44505+9416=53921$. When tokenized using ${\bm{T}}_{3}$, the sequence becomes $[44,505,+,9,416,=,53,921]$. This tokenized representation can then be reinterpreted as an arithmetic expression in base-$1000$.